OpenStudy (anonymous):

Whats the easy lazy person way to solve this? According to the Bohr model, the allowed energies of the hydrogen atom are given by the equation: E = (-21.7 x 10^-19)/n2 J. From what energy level must an electron fall to the n = 2 state to produce a line at 486.1 nm, the blue-green line in the visible hydrogen spectrum? A. 4 B. 3 C. 1 D. 5 E. 6

OpenStudy (anonymous):

Use the Planck relation (E = hv) to calculate the energy of a 486.1 nm photon. Now graph \[\Delta E = R_h \left( \frac{1}{2^2} - \frac{1}{n^2} \right)\]on your graphing calculator, or some calculator app on your desktop. You should be able to pick out the correct n for the E = dE you need by eye. Then do the calculation to be sure you've got it right. That way you don't need to do the calculation for all five possibilities.